# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 4.2 Filtered arcade reverse martingales
At the start of a farming season, a farmer might ask how the production of an agricultural produce should be kicked off so to achieve a target total production amount at the end of the season while minding the adverse weather forecast for the mid-season. If $X_{0}$ is the total amount produced by the end of the first day and $X_{1}$ is the total production at the end of the last day of the season, the farmer has a good idea about what $\pi^{X}$ is given his experience over many seasons. So, how should the farmer proceed? The answer is using a filtered arcade reverse martingale (FARM).
@@
Given an $X$-RAP $(I_{t}^{(1)})$, a one-arc FARM for $X\sim\pi^{X}$ on $[T_{0},T_{1}]$ is a stochastic process defined by $M_{t}^{-}=\mathbb{E}[X_{0}\,|\,\mathcal{G}_{t}^{I}]$, where $\mathcal{G}_{t}^{I}:=\sigma(I_{u}^{(1)}\,|\,t\leqslant u\leqslant T_{1})$ and $t\in[T_{0},T_{1}]$.
As opposed to a FAM, a FARM is not a martingale, and it is not necessarily interpolating between $X_{0}$ and $X_{1}$, either. Instead, it is a reverse martingale (also called backward martingale) with respect to the reverse filtration $(\mathcal{G}_{t}^{I})$, which follows from the tower property of conditional expectation: $\mathbb{E}[M_{s}^{-}\,|\,\mathcal{G}_{t}^{I}]=\mathbb{E}[\mathbb{E}[X_{0}\,|\,\mathcal{G}_{s}^{I}]\,|\,\mathcal{G}_{t}^{I}]=\mathbb{E}[X_{0}\,|\,\mathcal{G}_{t}^{I}]=M_{t}^{-}$ for any $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. A FARM only interpolates between $X_{0}$ and $X_{1}$ if $(X_{0},X_{1})$ is a discrete-time reverse martingale, i.e., $\mathbb{E}[X_{0}\,|\,X_{1}]=X_{1}$, since $M_{T_{0}}^{-}=X_{0}$ and $M_{T_{1}}^{-}=\mathbb{E}[X_{0}\,|\,X_{1}]$ by construction.
To write a FARM $M_{t}^{-}$ as a function of $I_{t}^{(1)}$ and $X_{1}$, one needs another relaxation of the Markov property as hinted at in Remark 3.11.
@@
Let $\mathcal{I}\subseteq\mathbb{R}^{+}$ be a real interval, and $\tau_{0}>\tau_{1}>\ldots\geqslant\inf\mathcal{I}$ such that $\tau=\{\tau_{0},\tau_{1},\ldots\}\subset\mathcal{I}$. The set $\tau$ may be finite, or contain countably many elements. A stochastic process $(Y_{t})_{t\in\mathcal{I}}$ is called reverse $\tau$-conditionally Markov if
@@
for any $(s,t)\in\mathcal{I}^{2}$ such that $s\leqslant t$, where $\mathcal{G}^{Y}_{t}=\sigma(Y_{u}\,|\,t\leqslant u\leqslant\sup\mathcal{I})$ and $\tau_{m(s)}=\min\limits_{i\in\mathbb{N}}\{\tau_{i}\,|\,\tau_{i}\geqslant t\}$.
@@
$M_{t}^{-}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{1},X_{0}=x}(I_{t}^{(1)})\,\mathrm{d}F^{X_{0}\,|\,X_{1}}(x)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{1},X_{0}=x}(I_{t}^{(1)})\,\mathrm{d}F^{X_{0}\,|\,X_{1}}(x)},$ (4.74)
for $t\in(T_{0},T_{1})$ if $(I_{t}^{(1)})$ is $\{T_{1},T_{0}\}$-reverse-Markov. In the previous section, we used standard RAPs to significantly simplify the SDEs that FAMs satisfy. For FARMs, the corresponding notion is the one of the reverse standard RAP.
@@
Following the same procedure used for FAMs to derive their SDEs, while replacing the conditional-Markov property by the reverse conditional-Markov property, and standard RAPs by reverse standard RAPs, one finds that there exists a reverse standard Brownian motion $(W_{t}^{-})_{t\in[T_{0},T_{1}]}$ adapted to the reverse filtration $(\mathcal{G}_{t}^{I})$ such that
$M_{t}^{-}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{0}\,|\,I_{s}^{(1)},X_{1}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}}{H_{1}(T_{1})H_{2}(s)-H_{1}(s)H_{2}(T_{1})}\,\mathrm{d}W_{s}^{-}.$ (4.75)
@@ 4.3 The n-arc FAM
Let $X=(X_{0},X_{1},\ldots,X_{n})$ be a random vector distributed according to a martingale coupling $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})$, where $\mu_{i}\in\mathcal{P}_{1}(\mathbb{R})$ for $i=0,\ldots,n$ and $\mu_{i}\leqslant_{\mathrm{cx}}\mu_{i+1}$ for $i=0,\ldots,n-1$. Given an $X$-RAP $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ on the partition $\{T_{0},T_{n}\}_{*}$, we aim at constructing an $(\mathcal{F}^{I}_{t})$-martingale $(M_{t})_{t\in[T_{0},T_{n}]}$ such that $M_{T_{i}}\overset{a.s.}{=}X_{i}$, for $i=0,\ldots,n$. A major difference in the $n$-arc case is that $(I_{t}^{(n)})$ must be $\{T_{0},T_{n}\}_{*}$-conditionally-Markov to guarantee the interpolation of $(M_{t})$.
@@
Given an $X$-RAP $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ that is $\{T_{0},T_{n}\}_{*}$-conditionally-Markov, an $n$-arc FAM for $X\sim\pi^{X}$ is a stochastic process of the form $M_{t}=\mathbb{E}[X_{n}\,|\,\mathcal{F}_{t}^{I}]=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,\\ X_{m(t)},I_{t}^{(n)}]$ where $m(t)=\max\{i\in\mathbb{N}\,|\,T_{i}\leqslant t\}$.
By the tower property of conditional expectation, $(M_{t})_{t\in[T_{0},T_{n}]}$ is a martingale with respect to $(\mathcal{F}_{t}^{I})_{t\in[T_{0},T_{n}]}$. Moreover, $M_{T_{i}}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{i}]=X_{i}$ since $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})$. Hence, $(M_{t})$ is an interpolating martingale with respect to $(\mathcal{F}_{t}^{I})$. We notice that, without the conditional-Markov property, $M_{T_{i}}=\mathbb{E}[X_{n}\,|\,\mathcal{F}_{T_{i}}^{I}]$ is not necessarily a.s. equal to $X_{i}$. As for the case $n=1$, in the remainder of this section, we assume that the driver $(D_{t})$ of $(I_{t}^{(n)})$ has a density function on $[T_{0},T_{n}]$.
@@
Let $M_{t}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM restricted to $(T_{0},T_{n})_{*}$. Then,
$M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\,\mathrm{d}F^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}}{\int_{\mathbb{R}}f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\,\mathrm{d}F^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}},$ (4.76)
where $F^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}$ is the distribution function of $X_{n}$ conditional on random vector $(X_{0},\ldots,\\ X_{m(t)})$. In particular:
@@
    If $(X_{0},\ldots,X_{m(t)},X_{n})$ is a continuous random variable, with density function $f^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}$, then
    $M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})f^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(y)\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})f^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(y)\,\mathrm{d}y}.$ (4.77)
@@
    $M_{t}=\frac{\sum\limits_{y}yf^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\mathbb{P}[X_{n}=y\,|\,X_{0},\ldots,X_{m(t)}]}{\sum\limits_{y}f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\mathbb{P}[X_{n}=y\,|\,X_{0},\ldots,X_{m(t)}]}.$ (4.78)
The proof follows along the lines of the one-arc case. Introducing the notation
$u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)=f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1},X_{n}=y}(I_{t}^{(n)}),$ (4.79)
it is often more convenient to write
$M_{t}=\frac{\int_{\mathbb{R}^{n-m(t)}}y\,u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1},X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1},y)}{\int_{\mathbb{R}^{n-m(t)}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1},X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1},y)}.$ (4.80)
We can apply Itô’s formula in the same fashion as in the one-arc case. An interesting pattern appears if the signal function of $(I_{t}^{(n)})$ satisfies $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$, that significantly simplifies the $n$-arc case.
@@
Let $M_{t}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM, where the signal function of $(I_{t}^{(n)})$ satisfies $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$. Then, $M_{t}=\mathbb{E}[X_{m(t)+1}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$.
@@
I also follows that
@@
$=\int_{\mathbb{R}^{n-m(t)-1}}\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1}]$
@@
Let $M_{t}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM. If the $X$-RAP $(I_{t}^{(n)})$ is standard and a semimartingale, satisfying the condition in Proposition 4.24, with driver covariance $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, such that $(t,x)\rightarrow\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}=x]$ is $C^{2}(((T_{0},T_{n})_{*}\setminus N)\times Im(I_{t}^{(n)}))$ where $N\subset(T_{0},T_{n})_{*}$ contains finitely many elements, then the following holds:
@@
    $M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}\mathrm{Var}[X_{m(s)+1}\,|\,X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s},$ (4.88)
@@
    $W_{t}=\int_{T_{0}}^{t}\frac{1}{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}\,\mathrm{d}N_{s},$ (4.89)
    $\,\mathrm{d}N_{t}$ $=\bigg{(}\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{m(t)+1})-H_{1}(T_{m(t)+1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{m(t)+1})H_{2}(t)-H_{1}(t)H_{2}(T_{m(t)+1})}$
@@
    and $Z_{t}=I_{t}^{(1)}-\sum\limits_{i=0}^{m(t)}g_{i}(t)X_{i}-\mu_{A}(t)$, and $J_{t}=\sum\limits_{i=0}^{m(t)}g_{i}^{\prime}(t)X_{i}+\mu_{A}^{\prime}(t)$.
@@
    The process $(W_{t})_{t\in[T_{0},T_{n}]}$ is an $(\mathcal{F}^{I}_{t})$-adapted standard Brownian motion on $[T_{0},T_{n}]$.
@@
$M_{t}$ $=X_{0}+\int_{T_{0}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}\mathrm{Var}[X_{m(s)+1}\,|\,X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s}$
$=X_{m(t)}+\int_{T_{m(t)}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}\mathrm{Var}[X_{m(s)+1}\,|\,X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s}.$ (4.91)

### v3

@@ 5.2 Filtered arcade reverse martingales
At the start of a farming season, a farmer might wonder how the production of his crops should be kicked off to meet the target yield by the end of the season, while taking into account the adverse weather forecast for the mid-season. If $X_{0}$ is the total amount produced by the end of the first day and $X_{1}$ is the total production at the end of the last day of the season, the farmer has a good idea about what $\pi^{X}$ is given his experience over many seasons. So, how should the farmer proceed? The answer is to use a filtered arcade reverse martingale (FARM).
@@
Given an $X$-RAP $(I_{t}^{(1)})$, a one-arc FARM for $X\sim\pi^{X}$ on $T_{0},T_{1}]$ is a stochastic process defined by $M_{t}^{-}=\mathbb{E}[X_{0}\mid\mathcal{G}_{t}^{I}]$, where $\mathcal{G}_{t}^{I}\mathrel{\mathop{\ordinarycolon}}=\sigma(I_{u}^{(1)}\mid t\leqslant u\leqslant T_{1})$ and $t\in[T_{0},T_{1}]$.
Unlike a FAM, a FARM is not a martingale, and is not necessarily interpolating between $X_{0}$ and $X_{1}$ either. Instead, it is a reverse martingale (also called backward martingale) with respect to the reverse filtration $({\mathcal{G}}_{t}^{I})$, which follows from the tower property of conditional expectation: $\mathbb{E}[M_{s}^{-}\mid{\mathcal{G}}_{t}^{I}]=\mathbb{E}[\mathbb{E}[X_{0}\mid{\mathcal{G}}_{s}^{I}]\mid{\mathcal{G}}_{t}^{I}]=\mathbb{E}[X_{0}\mid{\mathcal{G}}_{t}^{I}]=M_{t}^{-}$ for any $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. A FARM only interpolates between $X_{0}$ and $X_{1}$ if $(X_{0},X_{1})$ is a discrete-time reverse martingale, i.e., $\mathbb{E}[X_{0}\mid X_{1}]=X_{1}$, since $M_{T_{0}}^{-}=X_{0}$ and $M_{T_{1}}^{-}=\mathbb{E}[X_{0}\mid X_{1}]$ by construction.
To write a FARM $M_{t}^{-}$ as a function of $I_{t}^{(1)}$ and $X_{1}$, one needs another relaxation of the Markov property, as hinted at in Remark 4.25.
@@
Let $\mathcal{I}\subseteq\mathbb{R}^{+}$ be a real interval and $\tau_{0}>\tau_{1}>\ldots\geqslant\inf\mathcal{I}$ such that $\tau=\{\tau_{0},\tau_{1},\ldots\}\subset\mathcal{I}$. The set $\tau$ may be finite or contain countably many elements. A stochastic process $(Y_{t})_{t\in\mathcal{I}}$ is called reverse $\tau$-conditionally Markov if
@@
for any $(s,t)\in\mathcal{I}^{2}$ such that $s\leqslant t$, where $\mathcal{G}^{Y}_{t}=\sigma(Y_{u}\mid t\leqslant u\leqslant\sup\mathcal{I})$ and $\tau_{m(s)}=\min\limits_{i\in\mathbb{N}}\{\tau_{i}\mid\tau_{i}\geqslant t\}$.
@@
$M_{t}^{-}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\mid X_{1},X_{0}=x}(I_{t}^{(1)})\,\mathrm{d}F^{X_{0}\mid X_{1}}(x)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{1},X_{0}=x}(I_{t}^{(1)})\,\mathrm{d}F^{X_{0}\mid X_{1}}(x)},$ (5.65)
for $t\in(T_{0},T_{1})$ if $(I_{t}^{(1)})$ is reverse $\{T_{1},T_{0}\}$-conditionally Markov. In the previous section, we used standard RAPs to significantly simplify the SDEs that FAMs satisfy. For FARMs, the corresponding notion is the one of the reverse standard RAP.
@@
Following the same procedure used for FAMs to derive their SDEs, while replacing the conditionally-Markov property by the reverse conditionally-Markov property, and standard RAPs by reverse standard RAPs, one finds that there exists a reverse standard Brownian motion $(W_{t}^{-})_{t\in[T_{0},T_{1}]}$ adapted to the reverse filtration $(\mathcal{G}_{t}^{I})$ such that
$M_{t}^{-}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{0}\mid I_{s}^{(1)},X_{1}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}}{H_{1}(T_{1})H_{2}(s)-H_{1}(s)H_{2}(T_{1})}\,\mathrm{d}W_{s}^{-}.$ (5.66)
@@ 5.3 The n-arc FAM
Let $X=(X_{0},X_{1},\ldots,X_{n})$ be a random vector distributed according to a martingale coupling $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})$, where $\mu_{i}\in\mathcal{P}_{1}(\mathbb{R})$ for $i=0,\ldots,n$ and $\mu_{i}\leqslant_{\mathrm{cx}}\mu_{i+1}$ for $i=0,\ldots,n-1$. Given an $X$-RAP $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ on the partition $\{T_{0},T_{n}\}_{*}$, we aim at constructing an $(\mathcal{F}^{I}_{t})$-martingale $(M_{t})_{t\in[T_{0},T_{n}]}$ such that $M_{T_{i}}\overset{a.s.}{=}X_{i}$, for $i=0,\ldots,n$. A major difference in the $n$-arc case is that $(I_{t}^{(n)})$ must be $\{T_{0},T_{n}\}_{*}$-conditionally Markov to guarantee the interpolation of $(M_{t})$.
@@
Given an $X$-RAP $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ that is $\{T_{0},T_{n}\}_{*}$-conditionally Markov, an $n$-arc FAM for $X\sim\pi^{X}$ is a stochastic process of the form $M_{t}=\mathbb{E}[X_{n}\mid\mathcal{F}_{t}^{I}]=\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ where $m(t)=\max\{i\in\mathbb{N}\mid T_{i}\leqslant t\}$.
By the tower property of conditional expectation, $(M_{t})_{t\in[T_{0},T_{n}]}$ is a martingale with respect to $(\mathcal{F}_{t}^{I})_{t\in[T_{0},T_{n}]}$. Moreover, $M_{T_{i}}=\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{i}]=X_{i}$ since $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})$. Hence, $(M_{t})$ is an interpolating martingale with respect to $(\mathcal{F}_{t}^{I})$. We notice that, without the conditionally-Markov property, $M_{T_{i}}=\mathbb{E}[X_{n}\mid\mathcal{F}_{T_{i}}^{I}]$ is not necessarily a.s. equal to $X_{i}$. In other words, $X$ is not necessarily a martingale with respect to $(\mathcal{F}_{t}^{I})$. More precisely, if we consider the filtration $(\mathcal{F}_{t}^{X})_{t\in[T_{0},T_{n}]}$ generated by the martingale step process $(X_{t})_{t\in[T_{0},T_{n}]}$ equal to $X_{i}$ on $T_{i},T_{i+1})$ for $i=0,\ldots,n-1$ and to $X_{n}$ on $T_{n-1},T_{n}]$, we have that $\mathcal{F}_{t}^{X}\subseteq\mathcal{F}_{t}^{I}$ for all $t\in[T_{0},T_{n}]$ by definition of a RAP, but $(\mathcal{F}_{t}^{X})$ is not immersed [^4] We say that $\left(\mathcal{F}_{t}\right)$ is immersed in $\left(\mathcal{G}_{t}\right)$ under $\mathbb{P}$ if every $\left(\mathcal{F}_{t}\right)$-martingale is also a $\left(\mathcal{G}_{t}\right)$-martingale. The immersion property ensures that enlarging the filtration (i.e., gaining more information) does not alter the martingale property of processes adapted to the smaller filtration, meaning that the additional information in $\left(\mathcal{G}_{t}\right)$ does not allow for ”better prediction” of the future of $\left(\mathcal{F}_{t}\right)$-martingales. in $(\mathcal{F}_{t}^{I})$ under $\mathbb{P}$ in general. By restricting to RAPs that satisfy the $\{T_{0},T_{n}\}_{*}$-conditionally-Markov property, $(\mathcal{F}_{t}^{X})$ becomes immersed in $(\mathcal{F}_{t}^{I})$ under $\mathbb{P}$. Notice that for $n=1$, $(\mathcal{F}_{t}^{X})$ is immersed in $(\mathcal{F}_{t}^{I})$ under $\mathbb{P}$ in general. This is why imposing the conditionally-Markov property in the definition of a one-arc FAM is not necessary.
As for the case $n=1$, in the remainder of this section, we assume that the driver $(D_{t})$ of $(I_{t}^{(n)})$ has a density function on $T_{0},T_{n}]$.
@@
Let $M_{t}=\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM restricted to $(T_{0},T_{n})_{*}$. Then,
$M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\,\mathrm{d}F^{X_{n}\mid X_{0},\ldots,X_{m(t)}}}{\int_{\mathbb{R}}f^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\,\mathrm{d}F^{X_{n}\mid X_{0},\ldots,X_{m(t)}}},$ (5.67)
where $F^{X_{n}\mid X_{0},\ldots,X_{m(t)}}$ is the distribution function of $X_{n}$ conditional on random vector $(X_{0},\ldots,X_{m(t)})$. In particular:
@@
    If $(X_{0},\ldots,X_{m(t)},X_{n})$ is a continuous random variable, with density function $f^{X_{n}\mid X_{0},\ldots,X_{m(t)}}$, then
    $M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})f^{X_{n}\mid X_{0},\ldots,X_{m(t)}}(y)\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})f^{X_{n}\mid X_{0},\ldots,X_{m(t)}}(y)\,\mathrm{d}y}.$ (5.68)
@@
    $M_{t}=\frac{\sum\limits_{y}yf^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\mathbb{P}[X_{n}=y\mid X_{0},\ldots,X_{m(t)}]}{\sum\limits_{y}f^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\mathbb{P}[X_{n}=y\mid X_{0},\ldots,X_{m(t)}]}.$ (5.69)
The proof follows along the lines of the one-arc case. Introducing the notations
$u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)=f^{I_{t}^{(n)}\mid X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1},X_{n}=y}(I_{t}^{(n)})$ (5.70)
and $\tilde{F}\mathrel{\mathop{\ordinarycolon}}=F^{X_{m(t)+1},\ldots,X_{n-1},X_{n}\mid X_{0},\ldots,X_{m(t)}}$, it is often more convenient to write
$M_{t}=\frac{\int_{\mathbb{R}^{n-m(t)}}y\,u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n-1},y)}{\int_{\mathbb{R}^{n-m(t)}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n-1},y)}.$ (5.71)
We can apply Itô’s formula in the same fashion as in the one-arc case. An interesting pattern appears if the signal function of $(I_{t}^{(n)})$ satisfies $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$, which significantly simplifies the $n$-arc case.
@@
Let $M_{t}=\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM, where the signal function of $(I_{t}^{(n)})$ satisfies $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$. Then, $M_{t}=\mathbb{E}[X_{m(t)+1}\mid X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$.
@@
It also follows that
@@
$=\int_{\mathbb{R}^{n-m(t)-1}}\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1}]$
@@
Let $M_{t}=\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM. If the $X$-RAP $(I_{t}^{(n)})$ is standard and a semimartingale, satisfying the condition in Proposition 5.46, with driver covariance $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, such that $(t,x)\rightarrow\mathbb{E}[X_{n}\mid X_{0},\ldots,X_{m(t)},I_{t}^{(n)}=x]$ is $C^{2}(((T_{0},T_{n})_{*}\setminus N)\times Im(I_{t}^{(n)}))$ where $N\subset(T_{0},T_{n})_{*}$ contains finitely many elements, then the following holds:
@@
    $M_{t}$ $=X_{0}$
    $+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{m(s)+1}\mid X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s},$ (5.79)
@@
    $W_{t}=\int_{T_{0}}^{t}\frac{1}{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}}\,\mathrm{d}N_{s},$ (5.80)
    $\,\mathrm{d}N_{t}$
    $=\bigg(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{m(t)+1})-H_{1}(T_{m(t)+1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{m(t)+1})H_{2}(t)-H_{1}(t)H_{2}(T_{m(t)+1})}$
@@
    $Z_{t}=I_{t}^{(1)}-\sum\limits_{i=0}^{m(t)}g_{i}(t)X_{i}-\mu_{A}(t)$, and $J_{t}=\sum\limits_{i=0}^{m(t)}g_{i}^{\prime}(t)X_{i}+\mu_{A}^{\prime}(t)$.
@@
    The process $(W_{t})_{t\in[T_{0},T_{n}]}$ is an $(\mathcal{F}^{I}_{t})$-adapted standard Brownian motion on $T_{0},T_{n}]$.
@@
$M_{t}$
$=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{m(s)+1}\mid X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s}$
$=X_{m(t)}+\int_{T_{m(t)}}^{t}\frac{\mathrm{Var}[X_{m(s)+1}\mid X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s}.$ (5.82)